# Deepfake Generation and Detection: A Benchmark and Survey

### v3

@@ Abstract
Deepfake is a technology dedicated to creating highly realistic facial images and videos under specific conditions, which has significant application potential in fields such as entertainment, movie production, digital human creation, to name a few. With the advancements in deep learning, techniques primarily represented by Variational Autoencoders and Generative Adversarial Networks have achieved impressive generation results. More recently, the emergence of diffusion models with powerful generation capabilities has sparked a renewed wave of research. In addition to deepfake generation, corresponding detection technologies continuously evolve to regulate the potential misuse of deepfakes, such as for privacy invasion and phishing attacks. This survey comprehensively reviews the latest developments in deepfake generation and detection, summarizing and analyzing current state-of-the-arts in this rapidly evolving field. We first unify task definitions, comprehensively introduce datasets and metrics, and discuss developing technologies. Then, we discuss the development of several related sub-fields and focus on researching four representative deepfake fields: face swapping, face reenactment, talking face generation, and facial attribute editing, as well as forgery detection. Subsequently, we comprehensively benchmark representative methods on popular datasets for each field, fully evaluating the latest and influential published works. Finally, we analyze challenges and future research directions of the discussed fields.
Index Terms:
Deepfake Generation, Face Swapping, Face Reenactment, Talking Face Generation, Facial Attribute Editing, Forgery detection, Survey
@@ 1 Introduction
Artificial Intelligence Generated Content (AIGC) has garnered considerable attention [1] in academia and industry in recent years. Deepfake generation, as one of the important technologies in the generative domain, has gained significant attention due to its ability to create highly realistic facial media content. This technique has undergone a transition from traditional graphics-based methods to deep learning-based approaches. Early methods employed advanced Variational Autoencoder [2, 3, 4] (VAEs) and Generative Adversarial Networks [5, 6] (GANs) techniques, enabling seemingly realistic image generation, but its performance is still unsatisfactory that limits practical applications. Recently, the Diffusion [7, 8, 9] structure has greatly enhanced the generation capability of images/videos. Benefiting from this new wave of research, the deepfake technology allows us to see the potential values for the practical application, and it can generate indistinguishable contents against real ones, which has further attracted our attention and is widely applied in numerous fields [10], including entertainment, movie production, online live broadcasting, virtual meetings, new insights for privacy protection, etc.
Deepfake generation can generally be divided into four mainstream research fields: 1) Face swapping [11, 12, 13] is dedicated to executing identity exchanges between two person images; 2) Face reenactment [14, 15, 16] emphasizes transferring source movements and poses; 3) Talking face generation [17, 18, 19] focuses on achieving natural matching of mouth movements to textual content in character generation, and 4) Facial attribute editing [20, 21, 22] aims to modify specific facial attributes of the target image. The development of related foundational technologies has gradually shifted from single forward GAN models [23, 5] to multi-step diffusion models [24, 25, 7] with higher quality generation capabilities, and the generated content has also gradually transitioned from single-frame images to temporal video modeling [8]. In addition, NeRF [26, 27] has been frequently incorporated into modeling to improve multi-view consistency capabilities [28, 29].
@@
While enjoying the novelty and convenience of this technology, the unethical use of it raises concerns over the spread of privacy invasion, and dissemination of fake news, necessitating the development of effective forgery detection methods [30] in opposition, a.k.a., deepfake detection. From the earliest handcrafted feature-based methods [31, 32] to deep learning-based methods [33, 34], and the recent hybrid detection techniques [35], forgery detection has undergone substantial technological advancements along with the development of generative technologies. The data modality has also transitioned from the spatial and frequent domains [36, 37] to the more challenging temporal domain [38, 39]. Considering that current generative technologies have a higher level of interest, develop faster, and can generate indistinguishable content from reality [40], corresponding detection technologies need continuous evolution.
Overall, despite the significant progress made in both directions, they still exhibit challenging limitations in specific scenarios [41], mainly reflected in the visual perception authenticity and generative accuracy of models. This has attracted a large number of researchers to continue their efforts and has sparked thoughts on industrial applications. Existing survey works only focus on partial deepfake fields and lack discussions on new technologies [1, 40, 42], especially diffusion-based image/video generation methods, due to their disconnection from current technologies. The survey will comprehensively discuss these fields as well as related sub-fields, and cover tracking of the latest works.
Figure 2: Top: Illustration of different deepfake generation (Sec. 3.1) and detection (Sec. 3.2) tasks that are discussed in this survey. Bottom: Specific facial attribute modification of each task. Data from NVIDIA Keynote at COMPUTEX 2023 at 29:40.
@@
$\bullet$ Scope. This survey primarily focuses on mainstream face-related tasks, including face swapping, face reenactment, talking face generation, facial (multiple) attribute editing, and forgery detection. We also cover some related domain tasks in Sec. 2.4 and detail specific popular sub-tasks in Sec. 3.3. Considering the large number of articles (including published and preprints), we mainly include representative and attention-grabbing works. In addition, we compare this investigation with recent surveys. Sha et al[1] only discuss character generation while we cover a more comprehensive range of tasks. Compared to works [40, 42, 43], our study encompasses a broader range of technical models, particularly the more powerful diffusion-based methods. Additionally, we thoroughly discuss the related sub-fields of deepfake generation and detection.
$\bullet$ Survey Pipeline. Fig. 1 shows the pipeline of this survey. Sec. 2 provides the essential background knowledge that encompasses task-specific definitions, datasets, evaluation metrics, and exploration of related research areas. Sec. 3 presents technical discussions, examining four most popular deepfake tasks from the perspective of technological categorization and evolution. In addition, we meticulously classify and discuss forgery detection techniques, emphasizing the technological routes involved. Subsequently, Sec. 4 organizes and evaluates performance from various approaches for a thorough and fair performance comparison. Sec. 5 focuses on a critical review of the challenges that persist in existing techniques, outlining feasible future development directions. Finally, we encapsulate the entirety of the paper through a comprehensive summary in Sec. 6.
@@ 2 Background
In this section, we first introduce the conceptual definitions of the discussed mainstream fields. Fig. 2 illustrates intuitive objectives for each task and shows the distinctions among tasks in terms of manipulated facial components. Then, we review the developmental history of commonly used neural networks, highlighting several representative ones. Next, we summarize popular datasets, metrics, and loss functions. Finally, we comprehensively discuss several relevant domains.
@@ 2.1 Problem Definition
$\bullet$ Face Reenactment alters the facial movement of a target image $I_{t}$ without changing the identity and other facial attributes, under the conditions of driving image, video, head pose, etc.. This technology often involves the support of facial motion capture technology, such as facial tracking or prediction based on deep learning models [44].
@@ 2.2 Technological History and Roadmap
$\bullet$ Generative Framework. Variational Autoencoders (VAEs) [2, 3, 4], Generative Adversarial Networks (GANs) [23, 5, 6], and Diffusion [45, 24, 25] have played pivotal roles in the developmental history of generative models. 1) VAE [2] emerged in 2013, altering the relationship between latent features and linear mappings in autoencoder latent spaces. They introduced feature distributions like the Gaussian distribution and achieved the generation of new entities through interpolation. To enhance generation capabilities under specific conditions, CVAE [3] introduced conditional input. VQ-VAE [4] introduced the concept of Vector Quantization to improve the learning of latent representations. Subsequent models are continually advancing. 2) GANs [23] achieves high-quality generation through the adversarial training with an extra discriminator. Subsequently, research on GANs experienced a surge, and currently, the implementation based on GANs remains the mainstream approach for various deepfake tasks. CGAN [46] introduced conditional control variables to GANs. Pix2Pix [47] enhanced GAN performance in specific image translation tasks. StyleGAN [5] introduced the concept of style transfer to GANs. StyleGAN2 [6] further improved the quality and controllability of generated images. Additionally, some composite research combines GANs with VAE, e.g. CVAE-GAN [48]. 3) Diffusion modal [45] models the generation of data as a diffusion process. DDPM [24] gains widespread attention for its outstanding generative performance, especially when the model excels in handling large-scale, high-resolution images. LDM [25] is more flexible and powerful when it comes to modeling complex data distributions. In the field of video generation, Diffusion models play a crucial role [49, 50]. SVD [7] fine-tunes the base model using an image-to-video conversion task based on text-to-video models. AnimateDiff [8] attaches a newly initialized motion modeling module to a frozen text-to-image model and trains it in subsequent video clips to refine reasonable motion prior knowledge, thus achieving excellent generation results.
@@
$\bullet$ Discriminative Neural Network. Convolutional Neural Networks (CNNs) [51, 52, 53, 54, 55, 56] have played a pivotal role in the history of deep learning. LeNet [51], as the pioneer of CNNs, showcased the charm of machine learning. AlexNet [52] and ResNet [53] made deep learning feasible. Recently, ConvNeXt [54] has achieved excellent results surpassing those of Swin-Transformer [57]. The Transformer architecture initially proposed [58] in 2017. The core idea involves using self-attention mechanisms to capture dependencies between different positions in the input sequence, enabling global modeling of sequences. ViT [59] demonstrates that using Transformer in the field of computer vision can still achieve excellent performance and PVT [60] overcomes the challenges of adapting Transformer to various dense prediction tasks. In addition, Swin-Transformer [57] addresses the limitations of Transformer in handling high-resolution image processing tasks. Subsequently, Swin-Transformer V2 [61] further improves the model’s efficiency and the resolution of manageable inputs.
@@
$\bullet$ Neural Radiance Field (NeRF) is first introduced in 2020 [26], with its core idea revolving around the use of volume rendering and implicit neural fields to represent and reconstruct both geometric and illumination information of 3D scenes [27]. Compared to traditional 3D methods, it exhibits higher visual quality and is currently widely applied in tasks such as 3D geometry enhancement [62, 63], segmentation [64] and 6D pose estimation [65]. In addition, Some notable works [66, 67] combining NeRF as a supplement to 3D information and generation models are particularly prominent at present.
@@
TABLE I: Overview of commonly used datasets. Orange-marked ones are selected to evaluate different methods in Sec. 4.
Dataset Type Scale Highlight
Deepfake Generation LFW [68] Image 10K Facial images captured under various lighting conditions, poses, expressions, and occlusions at 250$\times$250 resolution.
CelebA [69] Image 200K The dataset includes over 200K facial images from more than 10K individuals, each with 40 attribute labels.
VGGFace [70] Image 2600K A super-large-scale facial dataset involving a staggering 26K participants, encompassing a total of 2600K facial images.
VoxCeleb1 [71] Video 100K voices A large scale audio-visual dataset of human speech, the audio includes noisy background interference.
CelebA-HQ [72] Image 30K A high-resolution facial dataset consisting of 30K face images, each with a resolution of 1024$\times$1024 resolution.
VGGFace2 [73] Image 3000K A large-scale facial dataset, expanded with greater diversity in terms of ethnicity and pose compared to VGGFace.
VoxCeleb2 [74] Video 2K hours A dataset that is five times larger in scale than VoxCeleb1 and improves racial diversity.
FFHQ [5] Image 70K The dataset with over 70K high-resolution (1024$\times$1024) facial images, showcasing diverse ethnicity, age, and backgrounds.
MEAD [75] Video 40 hours An emotional audiovisual dataset provides facial expression information during conversations with various emotional labels.
CelebV-HQ [76] Video 35K videos The dataset’s video clips have a resolution of more than 512$\times$512 resolution and are annotated with rich attribute labels.
Forgery Detection DeepfakeTIMIT [77] Audio-Video 640 videos The dataset is evenly divided into two versions: LQ (64$\times$64) and HQ (128$\times$128), with all videos using face swapping forgery.
FF++ [78] Video 6K videos Comprising 1K original videos and manipulated videos generated using five different forgery methods.
DFDCP [79] Audio-Video 5K videos The preliminary dataset for The Deepfake Detection Challenge includes two face-swapping methods.
DFD [80] Video 3K videos The dataset comprises 3K deepfake videos generated using five forgery methods.
Deeperforensics [81] Video 60K videos The videos feature faces with diverse skin tones, and rich environmental diversity was considered during the filming process.
DFDC [82] Audio-Video 128K videos The official dataset for The Deepfake Detection Challenge and contain a substantial amount of interference information.
Celeb-DF [83] Video 1K videos Comprising 408 genuine videos from diverse age groups, ethnicities, and genders, along with 795 DeepFake videos.
Celeb-DFv2 [83] Video 6K videos An expanded version of Celeb-DFv1, this dataset not only increases in quantity but also the diversity.
FakeAVCeleb [84] Audio-Video 20K videos A novel audio-visual multimodal deepfake detection dataset, deepfake videos generated using four forgery methods.

### v4

@@ Abstract
Deepfake is a technology dedicated to creating highly realistic facial images and videos under specific conditions, which has significant application potential in fields such as entertainment, movie production, and digital human creation, to name a few. With the advancements in deep learning, techniques primarily represented by Variational Autoencoders and Generative Adversarial Networks have achieved impressive generation results. More recently, the emergence of diffusion models with powerful generation capabilities has sparked a renewed wave of research. In addition to deepfake generation, corresponding detection technologies continuously evolve to regulate the potential misuse of deepfakes, such as privacy invasion and phishing attacks. This survey comprehensively reviews the latest developments in deepfake generation and detection, summarizing and analyzing current state-of-the-arts in this rapidly evolving field. First, we unify task definitions, comprehensively introduce datasets and metrics, and discuss developing technologies. Then, we discuss the development of several related sub-fields and focus on researching four representative deepfake fields: face swapping, face reenactment, talking face generation, and facial attribute editing, as well as forgery detection. Subsequently, we comprehensively benchmark representative methods on popular datasets for each field, fully evaluating the latest and most influential published works. Finally, we analyze the challenges and future research directions of the discussed fields. We closely follow the latest developments in this project.
Keywords:
Deepfake Generation Face Swapping Face Reenactment Talking Face Generation Facial Attribute Editing Forgery detection Survey
@@ 1 Introduction
Artificial Intelligence Generated Content (AIGC) garners considerable attention sha2023deep in academia and industry. Deepfake generation, as one of the important technologies in the generative domain, gains significant attention due to its ability to create highly realistic facial media content. This technique transitions from traditional graphics-based methods to deep learning-based approaches. Early methods employ advanced Variational Autoencoder kingma2014auto ; sohn2015learning ; van2017neural (VAE) and Generative Adversarial Network (GAN) karras2019style ; karras2020analyzing techniques, enabling seemingly realistic image generation, but their performance is still unsatisfactory, which limits practical applications. Recently, the diffusion structure blattmann2023stable ; guo2023animatediff ; liu2024towards has greatly enhanced the generation capability of images and videos. Benefiting from this new wave of research, deepfake technology demonstrates potential value for practical applications and can generate content indistinguishable from real ones, which has further attracted attention and is widely applied in numerous fields barattin2023attribute , including entertainment, movie production, online live broadcasting, and privacy protection, etc.
@@
Deepfake generation can generally be divided into four mainstream research fields: 1) Face swapping xu2022region ; ancilotto2023ximswap ; shiohara2023blendface is dedicated to executing identity exchanges between two person images; 2) Face reenactment koujan2020head2head ; hsu2022dual ; bounareli2023hyperreenact emphasizes transferring source movements and poses; 3) Talking face generation mir2023dit ; ma2023dreamtalk ; zhang2024dr2 focuses on achieving natural matching of mouth movements to textual content in character generation, and 4) Facial attribute editing sun2022anyface ; huang2023ia ; pang2023dpe aims to modify specific facial attributes of the target image. The development of related foundational technologies has gradually shifted from single forward GAN models goodfellow2014generative ; karras2019style to multi-step diffusion models ho2020denoising ; rombach2022high ; blattmann2023stable with higher quality generation capabilities, and the generated content has also gradually transitioned from single-frame images to temporal video modeling guo2023animatediff . In addition, NeRF mildenhall2020nerf ; gao2022nerf has been frequently incorporated into modeling to improve multi-view consistency capabilities jiang2022nerffaceediting ; zhang2022fdnerf .
While enjoying the novelty and convenience of this technology, the unethical use of it raises concerns over the spread of privacy invasion, and dissemination of fake news, necessitating the development of effective forgery detection methods tan2023learning in opposition, a.k.a., deepfake detection. From the earliest handcrafted feature-based methods he2019detection ; zhou2017two to deep learning-based methods yin2023dynamic ; yang2023masked , and the recent hybrid detection techniques ilyas2023avfakenet , forgery detection has undergone substantial technological advancements along with the development of generative technologies. The data modality has also transitioned from the spatial and frequent domains qian2020thinking ; li2021frequency to the more challenging temporal domain haliassos2022leveraging ; yang2023avoid . Considering that current generative technologies have a higher level of interest, develop faster, and can generate indistinguishable content from reality mirsky2021creation , corresponding detection technologies need continuous evolution.
Overall, despite the significant progress made in both directions, they still exhibit challenging limitations in specific scenarios vyas2024analysing , mainly reflected in the visual perception authenticity and generative accuracy of models. This has attracted a large number of researchers to continue their efforts and has sparked thoughts on industrial applications. Existing survey works only focus on partial deepfake fields and lack discussions on new technologies sha2023deep ; mirsky2021creation ; liu2023gan , especially diffusion-based image/video generation methods, due to their disconnection from current technologies. The survey will comprehensively discuss these fields as well as related sub-fields, and cover tracking of the latest works.
Figure 2: Top: Illustration of different deepfake generation (Sec. 3.1) and detection tasks (Sec. 3.2) that are discussed in this survey. Bottom: Specific facial attribute modification of each task. Data from NVIDIA Keynote at COMPUTEX 2023 at 29:40.
@@
$\bullet$ Scope. This survey primarily focuses on mainstream face-related tasks, including face swapping, face reenactment, talking face generation, facial (multiple) attribute editing, and forgery detection. We also cover some related domain tasks in Sec. 2.4 and detail specific popular sub-tasks in Sec. 3.3. Considering the large number of articles (including published and preprints), we mainly include representative and attention-grabbing works. In addition, we compare this investigation with recent surveys. Sha et alsha2023deep only discuss character generation while we cover a more comprehensive range of tasks. Compared to works mirsky2021creation ; liu2023gan ; melnik2024face , our study encompasses a broader range of technical models, particularly the more powerful diffusion-based methods. Additionally, we thoroughly discuss the related sub-fields of deepfake generation and detection.
$\bullet$ Survey Pipeline. Fig. 1 shows the pipeline of this survey. Sec. 2 provides the essential background knowledge that encompasses task-specific definitions, datasets, evaluation metrics, and exploration of related research areas. Sec. 3 presents technical discussions, examining the four most popular deepfake tasks from the perspective of technological categorization and evolution. In addition, we meticulously classify and discuss forgery detection techniques, emphasizing the technological routes involved. Subsequently, Sec. 4 organizes and evaluates performance from various approaches for a thorough and fair performance comparison. Sec. 5 focuses on a critical review of the challenges that persist in existing techniques, outlining feasible future development directions. Finally, we encapsulate the entirety of the paper through a comprehensive summary in Sec. 6.
@@ 2 Background
In this section, we first introduce the conceptual definitions of the discussed mainstream fields. Fig. 2 illustrates the intuitive objectives for each task and shows the distinctions among tasks in terms of manipulated facial components. Then, we review the developmental history of commonly used neural networks, highlighting several representative ones. Next, we summarize popular datasets, metrics, and loss functions. Finally, we comprehensively discuss several relevant domains.
@@ 2.1 Problem Definition
$\bullet$ Face Reenactment alters the facial movement of a target image $I_{t}$ without changing the identity and other facial attributes, under the conditions of driving image, video, head pose, etc.. This technology often involves the support of facial motion capture technology, such as facial tracking or prediction based on deep learning models zheng2021efficient .
@@ 2.2 Technological History and Roadmap
$\bullet$ Generative Framework. Variational Autoencoders (VAEs) kingma2014auto ; sohn2015learning ; van2017neural , Generative Adversarial Networks (GANs) goodfellow2014generative ; karras2019style ; karras2020analyzing , and Diffusion sohl2015deep ; ho2020denoising ; rombach2022high have played pivotal roles in the developmental history of generative models.
1) VAE kingma2014auto emerges in 2013, altering the relationship between latent features and linear mappings in autoencoder latent spaces. It introduces feature distributions like the Gaussian distribution and achieves the generation of new entities through interpolation. To enhance generation capabilities under specific conditions, CVAE sohn2015learning introduces conditional input. VQ-VAE van2017neural introduces the concept of vector quantization to improve the learning of latent representations. Subsequent models continually advance.
2) GANs goodfellow2014generative achieve high-quality generation through adversarial training with an extra discriminator. Subsequently, research on GANs experiences a surge, and currently, implementations based on GANs remain the mainstream approach for various deepfake tasks. CGAN mirza2014conditional introduces conditional control variables to GANs. Pix2Pix isola2017image enhances GAN performance in specific image translation tasks. StyleGAN karras2019style introduces the concept of style transfer to GANs. StyleGAN2 karras2020analyzing further improves the quality and controllability of generated images. Additionally, some composite research combines GANs with VAEs, e.g., CVAE-GAN bao2017cvae .
3) Diffusion models sohl2015deep model the generation of data as a diffusion process. DDPM ho2020denoising gains widespread attention for its outstanding generative performance, especially when the model excels in handling large-scale, high-resolution images. LDM rombach2022high is more flexible and powerful when it comes to modeling complex data distributions. In the field of video generation, diffusion models play a crucial role ma2023follow ; liu2023residual ; gupta2023photorealistic . SVD blattmann2023stable fine-tunes the base model using an image-to-video conversion task based on text-to-video models. AnimateDiff guo2023animatediff attaches a newly initialized motion modeling module to a frozen text-to-image model and trains it on subsequent video clips to refine reasonable motion prior knowledge, thus achieving excellent generation results.
Figure 4: Works summarization on different directions per year. Data is obtained on 2024/05/13.
Table 1: Overview of commonly used datasets. Orange-marked ones are selected to evaluate different methods in Sec. 4.
Dataset Type Scale Highlight
Deepfake Generation LFW huang2008labeled Image 10K Facial images captured under various lighting conditions, poses, expressions, and occlusions at 250$\times$250 resolution.
CelebA liu2015deep Image 200K The dataset includes over 200K facial images from more than 10K individuals, each with 40 attribute labels.
VGGFace parkhi2015deep Image 2600K A super-large-scale facial dataset involving a staggering 26K participants, encompassing a total of 2600K facial images.
VoxCeleb1 nagrani2017voxceleb Video 100K voices A large scale audio-visual dataset of human speech, the audio includes noisy background interference.
CelebA-HQ karras2018progressive Image 30K A high-resolution facial dataset consisting of 30K face images, each with a resolution of 1024$\times$1024 resolution.
VGGFace2 cao2018vggface2 Image 3000K A large-scale facial dataset, expanded with greater diversity in terms of ethnicity and pose compared to VGGFace.
VoxCeleb2 chung2018voxceleb2 Video 2K hours A dataset that is five times larger in scale than VoxCeleb1 and improves racial diversity.
FFHQ karras2019style Image 70K The dataset with over 70K high-resolution (1024$\times$1024) facial images, showcasing diverse ethnicity, age, and backgrounds.
MEAD wang2020mead Video 40 hours An emotional audiovisual dataset provides facial expression information during conversations with various emotional labels.
CelebV-HQ zhu2022celebv Video 35K videos The dataset’s video clips have a resolution of more than 512$\times$512 resolution and are annotated with rich attribute labels.
Forgery Detection DeepfakeTIMIT korshunov2018deepfakes Audio-Video 640 videos The dataset is evenly divided into two versions: LQ (64$\times$64) and HQ (128$\times$128), with all videos using face swapping forgery.
FF++ rossler2019faceforensics++ Video 6K videos Comprising 1K original videos and manipulated videos generated using five different forgery methods.
DFDCP dolhansky2019deepfake Audio-Video 5K videos The preliminary dataset for The Deepfake Detection Challenge includes two face-swapping methods.
DFD DfD Video 3K videos The dataset comprises 3K deepfake videos generated using five forgery methods.
Deeperforensics jiang2020deeperforensics Video 60K videos The videos feature faces with diverse skin tones, and rich environmental diversity was considered during the filming process.
DFDC dolhansky2020deepfake Audio-Video 128K videos The official dataset for The Deepfake Detection Challenge and contain a substantial amount of interference information.
Celeb-DF li2020celeb Video 1K videos Comprising 408 genuine videos from diverse age groups, ethnicities, and genders, along with 795 DeepFake videos.
Celeb-DFv2 li2020celeb Video 6K videos An expanded version of Celeb-DFv1, this dataset not only increases in quantity but also the diversity.
FakeAVCeleb khalid2021fakeavceleb Audio-Video 20K videos A novel audio-visual multimodal deepfake detection dataset, deepfake videos generated using four forgery methods.
@@
$\bullet$ Discriminative Neural Network. Convolutional Neural Networks (CNNs) lecun1998gradient ; krizhevsky2012imagenet ; he2016deep ; liu2022convnet ; emo ; eatformer have played a pivotal role in the history of deep learning. LeNet lecun1998gradient , as the pioneer of CNNs, showcased the charm of machine learning. AlexNet krizhevsky2012imagenet and ResNet he2016deep made deep learning feasible. Recently, ConvNeXt liu2022convnet has achieved excellent results surpassing those of Swin-Transformer liu2021swin . The Transformer architecture initially proposed vaswani2017attention in 2017. The core idea involves using self-attention mechanisms to capture dependencies between different positions in the input sequence, enabling global modeling of sequences. ViT dosovitskiy2020image demonstrates that using Transformer in the field of computer vision can still achieve excellent performance and PVT wang2021pyramid overcomes the challenges of adapting Transformer to various dense prediction tasks. In addition, Swin-Transformer liu2021swin addresses the limitations of Transformer in handling high-resolution image processing tasks. Subsequently, Swin-Transformer V2 liu2022swin further improves the model’s efficiency and the resolution of manageable inputs.
@@
$\bullet$ Neural Radiance Field (NeRF) is first introduced in 2020 mildenhall2020nerf , with its core idea revolving around the use of volume rendering and implicit neural fields to represent and reconstruct both geometric and illumination information of 3D scenes gao2022nerf . Compared to traditional 3D methods, it exhibits higher visual quality and is currently widely applied in tasks such as 3D geometry enhancement sharma2023neural ; irshad2023neo , segmentation liu2023instance and 6D pose estimation min2023neurocs . In addition, Some notable works hwang2023faceclipnerf ; yang2023urbangiraffe combining NeRF as a supplement to 3D information and generation models are particularly prominent at present.